OpenStudy (anonymous):

EASY QUESTION- MEDAL when doing basic algebraic expressions, how do you multiply variables? i.e. 3y * 5y i am having the worst brain fart ever right now it is not even funny

OpenStudy (ribhu):

first multiply 3*5

OpenStudy (anonymous):

so its just 15y?

OpenStudy (ribhu):

then multiply y*y = (y^1)(y^1)

OpenStudy (anonymous):

Fairly certain it's \[3y * 5y = 15y^{2}\]

OpenStudy (anonymous):

Multiply the whole numbers together, then the variables.

OpenStudy (anonymous):

@Nibby Please do not give answers, it's against OpenStudy's Code-of-Conduct. We are here to guide the asker to the answer, so we need his/her participation. http://openstudy.com/code-of-conduct

OpenStudy (ribhu):

then y^(1+!) = y^2

OpenStudy (anonymous):

Because: \[5 * 3 = 15\] and: \[y * y = y^{2}\]

OpenStudy (anonymous):

I was getting to it @iGreen.

OpenStudy (ribhu):

then the answer becomes 15 y^2

OpenStudy (anonymous):

so 3*5 = 15 and y*y = y^2 so 15y^2?

OpenStudy (anonymous):

@iGreen. LOL your doing it also -_-

OpenStudy (anonymous):

ok, thank you all its been a long week haha

OpenStudy (anonymous):

@Nibby \(\color{blue}{\text{Originally Posted by}}\) @iGreen. @Nibby Please do not give answers, it's against OpenStudy's Code-of-Conduct. We are here to guide the asker to the answer, so we need his/her PARTICIPATION. http://openstudy.com/code-of-conduct \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

I didn't do anything.. @Shikatanori all I said was: \(\color{blue}{\text{Originally Posted by}}\) @iGreen. Multiply the whole numbers together, then the variables. \(\color{blue}{\text{End of Quote}}\) She gave the answer..

OpenStudy (anonymous):

oh god please dont start another one of these fueds on my post

OpenStudy (anonymous):

I didnt say here ? :3

OpenStudy (ribhu):

did u get it? @kah.x

OpenStudy (anonymous):

Why did you restate the CoC response? @iGreen. I explained it soon after, apologies.

OpenStudy (anonymous):

Yes, but you weren't supposed to give it without the asker's participation.

OpenStudy (anonymous):

sTaHpPpPpP

OpenStudy (anonymous):

Just Add the variables exponent @kah.x

OpenStudy (anonymous):

then multiply the numerical value